# GLOBALIZATION? TRADE WAR? A COUNTERBALANCE PERSPECTIVE

### v4

@@ 6 Empirical Study
This empirical analysis calculates $\pi$ and its derivatives with respect to $P_{ii}$ and $P_{ji}$ for the years 2000 to 2019. The bilateral trade data come from the ComTrade database of the United Nations (2021). We use the import and export data not for consumer goods from the World Bank’s WITS Database (2021). Due to the rationality of importers and exporters, the data have largely reflected political, territorial, ideological, cultural, war, national security and other geopolitical considerations, as well as transaction and shipping costs, resource endowment, industrial distribution, and location advantages. For comparison, we focus on the years 2000, before China entered the WTO in 2001, and 2017, before the beginning of the China-USA trade war in 2018. During this period, there were a subprime mortgage crisis between 2007 and 2010 and a European debt crisis between 2009 and 2011. All these events heavily impacted relevant countries. Though all economies in the data are used in the calculations, we only present the results for China, Russia, and the G7 countries (Canada, France, Germany, Italy, Japan, the United Kingdom, and the USA). Their ISO-3 country codes are CHN, RUS, CAN, FRA, DEU, ITA, JPN, GBR, and USA, respectively. We use $\lambda_{ji}=\sqrt{\pi_{j}g_{j}}/\sqrt{\pi_{i}g_{i}}$ in the computations. The results are in the following tables and figures.
@@
First, Figure IV plots $\pi_{i}$ from 2000 to 2019. China and Russia gained, and Canada remained almost unchanged while six others lost. Specifically, the USA’s $\pi_{i}$ steadily grew before 2003; it then sharply dropped till 2011 and slowly recovered. Overall, it dropped by 12.7%. China’s $\pi_{i}$ sharply increased during the two crises but almost leveled off elsewhere. Canada and Russia showed their resilience in $\pi$ because of rising prices of commodities in these turmoiling years. The four Western European countries suffered significantly from the debt crisis, and their $\pi_{i}$ continued falling till 2015; manufacturing-focused Germany started falling in 2000 while three others from 2008. Japan’s $\pi_{i}$ plummeted nearly half in the first eight years but finally stopped this momentum between 2007 and 2014. It suffered a domino effect from 2000 to 2007 when China and South Korea challenged its automobile and electric industries. Beyond these nine countries, the others overall became more competitive with their total competitiveness improved by 22.7%, as seen in Figure VI. Many southern asian-, southeast asian-, and middle east countries had tripled their competitiveness including, but not limited to, India, Indonesia, Iraq, United Arab Emirates, and Vietnam.
Also, the Matthew effect was clearly demonstrated. For example, the USA and China’s production ratio was $1.344$ in 2018, but their $\pi_{i}$ ratio was $3.070$. The effect can also be evidenced by Figure VI, in which $\pi_{i}$ has a strong positive correlation with $g_{i}^{2}$ after controlling $g_{i}$. The partial correlation, however, declined from .757.757.757.757 to .527.527.527.527 during the study period when the trade system shifted from less to more multipolar.
@@
[…]
@@
Figure VII plots the posterior bargaining power from 2000 to 2019. It has similar but smoother patterns as Figure IV. Due to rapid production growth, China’s bargaining power steadily rose, including in 2005 and 2015, when it had troughs in $\pi_{i}$. Also, its $\beta_{i}>\pi_{i}$, whereas the USA’s $\beta_{i}<\pi_{i}$ in each year, showing $\beta_{i}$’s mitigation on the Matthew effect in $\pi_{i}$. Additionally, Russia’s bargaining power sharply dropped after the Russo-Georgian War in 2008 and its annexation of Crimea in 2014. Figure VIII compares the bargaining power versus the USA. The four western European countries’ bargaining power against the USA slightly increased after introducing the Eurozone in 1999 and modestly decreased after 2008. On average, they lost 15.7% of their bargaining power to the USA. Besides, China and Russia improved their relative power multiple times in the first 15 years and remained constant after that. Meanwhile, Japan lost about 50%. Modestly immune to the crises, Canada’s relative bargaining strength sharply rose during the crises and quickly returned to its original level due to its strong alignment with the USA.
@@
Next, Table II lists the derivatives of $\log\pi_{i}$ with respect to $\log P_{ji}$ for 2000 and 2017, calculated from Eq. (9) and (12). If the USA had cut 1% of its imports from China in 2017, for example, its $\pi_{i}$ would have increased by .04889%. While the USA may have had reasons to launch trade wars against all eight others, its top target was China in 2017 and Japan in 2000, according to the numerators and denominators of the last row. As seen from the last column, China and Japan would have been the two largest beneficiaries if they had further collaborated with the USA in 2017 and 2000, respectively. However, comparing the numerators and denominators in the column, the benefits were considerably reduced from 2000 to 2017. Also, starting trade friction from any of these eight countries against the USA would be a loss to the country. Additionally, any other country’s trade war against Russia would be nearly fruitless to increase its $\pi_{i}$, since both numerators and denominators are small in the eighth column. Also, the numbers symmetric with the diagonal have opposite signs, verifying Corollary 2.
@@
Table III lists the derivatives of $\log\pi$ with respect to $\log P_{ii}$ for 2017 and 2000, calculated from Eq. (22), (23), and (12). From the diagonal, globalization significantly helped China, Germany, Italy, Japan, and Russia. It was a good choice for Canada in 2000 but a bad one in 2017. The effects on France were insignificant because of its small numerator and denominator. The USA and the United Kingdom would be anti-globalization advocates; this explains recent events such as the Trump administration’s withdrawal from international organizations and the United Kingdom’s Brexit. Though business interest groups overwhelmingly profited from the gigantic international markets and low-priced raw materials, growing low-income Americans and immigrants demanded more inexpensive goods from China, uplifting trade deficits and national debts. Also, China was the largest beneficiary of globalization in 2017 with a 2.324% hike in $\pi_{i}$ for a 1% decrease in $P_{ii}$, so its Belt and Road Initiative (BRI) would have a positive impulse on its $\pi_{i}$. Russia was the largest beneficiary of globalization in 2000. The side effects on third parties are off the diagonal. For example, the USA’s further globalization benefited all eight other countries, and so its protectionism would meet serious resistance from its partners. China was the largest beneficiary from the USA’s globalization in both years. Canada suffered the least with the USA’s protectionism in 2017; its further globalization would have slightly profited the USA, according to the first row. The last column implies that further globalization from seven other countries would hurt the USA’s $\pi_{i}$; Japan and China topped the list in 2000 and 2017, respectively. Their hurt declined by 70.2% and ascended by 380%, respectively, in the 17 years. In addition, the second row helps to explain the USA and its western allies’ objection to China’s BRI-alike initiatives, intensified by $4.74$ times from 2000 to 2017 according to the ratios of their numerators and denominators.
@@
Finally, we evaluate the side effects on the third parties due to the USA’s trade war against China in 2018 and 2019. Table IV is calculated from Eq. (9), (11), and (12). When the USA reduced its imports from China by 1% in 2018 and 2019, China’s $\pi_{j}$ suffered a significant loss of .1058% and .0937%, respectively, while the USA’s $\pi_{i}$ earned a non-significant gain of .0540% and .0460%, respectively. Canada was a free rider, and the six other countries were slightly negatively affected in 2018 and 2019. Therefore, these countries would likely take no formal action to prevent the war. Also, Canada or Russia would be the best ally with the USA or China, respectively, in this conflict. The second best would be the United Kingdom for the USA and Japan for China, respectively. From 2018 to 2019, the effects were slightly mitigated for all these countries. Probably realizing the mutually destructive consequences of the trade war, China and the USA agreed on a new trade deal in early 2020 (cf, US Trade Representative, 2020), concluding the first phase of a long-term battle.
@@
The cooperative solution $\lambda_{ji}^{*}$ could mitigate the conflict and prevent further escalation. At the inception of this conflict in 2018,
@@
For any positive $\mathrm{d}P_{ji}$, therefore, any $\lambda_{ji}$ between $\pi_{j}/\pi_{i}$ and $g_{j}/g_{i}$ would not only make $\frac{\mathrm{d}\log\pi_{i}}{\mathrm{d}\log P_{ji}}<0$ but also create further trade deficits for the USA (see Table V). Thus, an economic conflict with a negative $\mathrm{d}P_{ji}$ was possible and, in this lose-lose conflict, the Nash bargaining solution was
@@
for 2018. For each dollar’s decrease in exports from China to the USA, $\mathrm{d}P_{ji}=-\$1/g_{j}$. According to the bargaining solution, China would have changed its imports from the USA by $\mathrm{d}P_{ij}=\lambda_{ji}^{*}\mathrm{d}P_{ji}$ of the USA’s production, which is .$66164.49233*(-1/g_{j})*g_{i}=-\$.$66164. So if this resolution had been applied, the USA would have earned \$.33836=(-.66164)-(-1) surplus for each decreased dollar of imports from China. Also, by Theorem  2, the USA’s $\pi_{i}$ would have increased since $\lambda^{*}_{ji}>\pi_{j}/\pi_{i}$ and $\mathrm{d}P_{ji}<0$. To maintain its competitive advantage $\pi_{i}$ and reduce national debt by earning a trade surplus in the bilateral friction, by the impossibility trilemma in Theorem 4 , the USA had to sacrifice its exports and shrink its imports — in either case, $\mathrm{d}P_{ij}<0$.
@@ 7 Discussions
Economic globalization and trade wars have emerged as broad public interests because of global competition among countries, large-scale restructuring of labor markets, and international distribution of added value. We argue that import and export data have already reflected these issues. Therefore, we attempt to extract a country’s competitive strength and bargaining power from the data, identify its best strategies for dealing with trade conflicts and cooperation, and determine if it should move more or less toward globalization. The employed network- and game-theoretical methodologies focus on contemporaneously interactive relationships among economies and are free of parameters and pre-set utility functions. The network game approaches could be applied to other bilateral flow or dyadic data (such as traffic, social networks, and global payment systems) and similar conflicts (such as currency wars and technology wars).
The forgoing account, however, passes over several vital factors that deserve further investigation. First, counterbalance disequilibrium may exist when trades are notably imbalanced. This could happen in many ways. For example, when country $i$ uses borrowed money to consume $j$’s exports, its consumption still has power over $j$’s production. However, the power comes not only from $i$, but also from the money lender, which has some control over $i$’s consumption. If the lender happens to be $j$, then $\pi_{i}$ is overestimated and $\pi_{j}$ underestimated. Secondly, unlike an economic superpower, an emerging economy may focus on its absolute growth of exports instead of competitive advantages. During China’s primitive capital accumulation in the 1980s and 1990s, for example, earning extra US dollars was necessary to finance it industrial modernization. Thus, maximizing $\pi_{i}$ may be a low priority for an emerging economy. The time dimension, absent from the current research, is required to formulate growth. Thirdly, we have yet to include economic globalization’s capital, information, and technology sides. They interplay with goods and services, but their data are not as easily accessible as goods and services. Also, several data issues may cast doubt on the results of the empirical study. There are many missing values in the service data, and country $i$’s reported imports from $j$ are often inconsistent with $j$’s reported exports to $i$. Next, the integration of regional economies merits further attention. Theorem 2 could be used to find the suitable candidates for the integration. If country $i$ belongs to several RTAs, however, $\Delta P_{ii}$ may not be proportionally offset by other countries in Theorem 6. Lastly, one could add other variables to the demand-sided counterbalance Eq. (3), especially those from the supply side, such as labor mitigation, which may fine-tune the fractions in $P$. Any factors directly impacting $i$’s production could be counted in the $i$th row of $P$. Though consumption may still dominate the fractions, not all productions are driven by consumption.
One could extend the framework in several directions. First, between a bilateral trade war and anti-globalization, a country could target a select group of partners without directly affecting others. For example, as Table II indicates, the USA could simultaneously raise tariffs on imports from China, Germany, and Japan. Secondly, though the dynamical system of Eq. (9)–(11) or Eq. (22)–(23) assumes a fixed $P$, we could include a time dimension to $P$. With the time horizon, we could discuss a trade war’s effects on inflation, $\pi_{i}$’s dependence on its historical values, and the improvement of trade deficits. The time dimension could also improve the forecast capacity of the model. Intertemporal delay often exists when production happens in one year but consumption in the following years. If the production takes two years, then the previous year’s imports of intermediate goods could be included in the current year’s completed production. Next, we could let $\lambda_{ji}$ depend on other determinants besides the competitiveness and production ratios. An industrial or geopolitical analysis could shed light on the choice. For example, as Japan and South Korea rivaled for automobile and electronic exports in 2000, $\pi_{j}/\pi_{i}$ was proper for $\lambda_{ji}$. In the same year, no industry could be a battlefield for China and the USA, so $g_{j}/g_{i}$ seemed a natural choice for $\lambda_{ji}$. Also, in the literature (e.g., Allen, Arkolakis, and Takahashi, 2020; Isard, 1954), the distance between trade partners is a vital determinant of their trade flows. Thus, one may discount the $i$th row of $P$ by the distances between country $i$ and its trade partners, partially removing location advantage from $\pi$. Econometric analyses could also bring additional insights. In the linear regression of Eq. (8), for example, if the estimated residual $\hat{\epsilon}_{i}$ is larger or less than 00, we may say $\pi_{i}$ is over- or under-valued, respectively. There are other regression forms and $(\pi_{i},g_{i})$ could be panel data. Inclusion of additional regressors, such as debt to GDP ratio, to the regression helps better understand $\pi$ by explanatory data. Finally, anticipating the side effects of $i$’s trade war against $j$, country $j$ could seek cooperation with third parties, which would substantially suffer from the conflict. However, adding more countries to both sides may result in a world war of trade, just like the Cold War between 1947 and 1991.
Reference
@@
We multiply $\vec{1}_{n}$ on both right sides of the equation $\rho P=c\rho$ to get $\rho P\vec{1}_{n}=c\rho\vec{1}_{n}.$ Using $P\vec{1}_{n}=\vec{1}_{n}$ and $\rho\vec{1}_{n}=1$, we then obtain $\rho\vec{1}_{n}=c.$ Thus, $c=1$. Also, by $\rho P=c\rho=\rho$ and the uniqueness of $\pi$, we have $\rho=\pi$.
A2. Proof of Theorem 2
We apply matrix calculus with the restriction of Eq. (3) to prove the theorem. When making a small shock or perturbation $\Delta P$ to $P$ in Eq. (3), the new authority distribution $\pi+\Delta\pi$ satisfies the counterbalance equation of
@@
subject to $\Delta P\vec{1}_{n}=\vec{0}_{n}$ and $\Delta\pi\vec{1}_{n}=0$. After subtracting $\pi=\pi P$ from Eq. (A.1), we get $\Delta\pi[I_{n}-P]=\pi\Delta P+\Delta\pi\Delta P$ and its first-order approximation is

### v5

@@ 6 Empirical Study
In this empirical study, we apply our theoretical results to real-world trade data spanning from 2000 to 2019. This analysis provides quantitative evidence to support our previous arguments, making the theoretical insights more tangible and relevant for policymakers. The empirical analysis reveals trends in national competitiveness and bargaining power over time, highlighting how different countries have gained or lost. This provides a historical perspective on the impacts of globalization and trade wars.
Between 2000 and 2019, the world economy underwent significant transformations and challenges, albeit with varying degrees of success across different regions. This period witnessed a substantial increase in global trade volumes, driven by rapid technological advancements, the expansion of international trade agreements, and the rise of emerging markets. The integration of China into the WTO in 2001 was a pivotal moment, leading to a surge in global trade activities. The rise of emerging markets, particularly China and India, played a crucial role in global economic growth but also led to a trade war between China and the USA. The 2008 financial crisis and the subsequent European debt crisis were major turning points, causing widespread economic downturns and leading to substantial policy interventions by governments and central banks worldwide.
As trade volumes increased with globalization, nations demanded more of the world’s reserve currency, primarily through trade surpluses. Consequently, the reserve currency appreciates, decreasing the currency issuer’s exports and increasing trade deficits. The issuer may have three options: encouraging local currency settlement in international transactions or reducing its currency’s usage as the world’s reserve currency; ignoring the trade deficits and focusing on national competitiveness while still increasing exports; or rolling back globalization worldwide. For the USA, reshoring its manufacturing sectors would enhance its competitiveness and grow exports; however, trade deficit may still be an issue according to the impossibility trilemma. For China, significant growth is necessary to shrink the competitiveness gap with the USA, according to the Matthew effect in $\pi$.
This empirical analysis calculates $\pi$ and its derivatives with respect to $P_{ii}$ and $P_{ji}$. The bilateral trade data are sourced from the United Nations ComTrade database (2021). We use import and export data for non-consumer goods from the World Bank’s WITS Database (2021). For comparison, we focus on the years 2000, before China entered the WTO in 2001, and 2017, before the beginning of the China-USA trade war in 2018. Though all economies in the data are used in the calculations, we only present the results for China, Russia, and the G7 countries (Canada, France, Germany, Italy, Japan, the UK, and the USA). Their ISO-3 country codes are CHN, RUS, CAN, FRA, DEU, ITA, JPN, GBR, and USA, respectively. We use $\lambda_{ji}=\sqrt{\pi_{j}g_{j}}/\sqrt{\pi_{i}g_{i}}$ in the computations. The results are in the following tables and figures.
@@
Figure IV illustrates the competitiveness trends from 2000 to 2019. During this period, China and Russia experienced significant gains, while Canada remained relatively stable, and the other six countries saw declines. Notably, the USA’s competitiveness grew steadily until 2003, followed by a sharp decline until 2011, and a gradual recovery thereafter, resulting in an overall decrease of 12.7%. China’s competitiveness surged during the two crises but plateaued otherwise. Canada and Russia demonstrated resilience due to rising commodity prices during transformative years. The four Western European countries were significantly impacted by the debt crisis, with their competitiveness continuing to fall until 2015. Germany, focused on manufacturing, began its decline in 2000, while the other three countries started in 2008. Japan’s competitiveness nearly halved in the first eight years but stabilized between 2007 and 2014, suffering a domino effect from 2000 to 2007 as China and South Korea challenged its automobile and electric industries. Beyond these nine countries, others collectively improved their competitiveness by 22.7%, as depicted in Figure VI. Many Southern Asian, Southeast Asian, and Middle East countries, including India, Indonesia, Iraq, United Arab Emirates, and Vietnam, tripled their competitiveness.
The Matthew effect is clearly demonstrated. For instance, the USA and China’s production ratio was $1.344$ in 2018, but their competitiveness ratio was $3.070$. This effect is further evidenced by Figure VI, which shows a strong positive correlation between $\pi_{i}$ and $g_{i}^{2}$ after controlling $g_{i}$. The partial correlation, however, declined from .757.757.757.757 to .527.527.527.527 during the study period when the trade system shifted from less to more multipolar.
@@
A straightforward implication of Eq. (4) is country $j$’s direct contribution to country $i$’s competitiveness in percentage, i.e., $\pi_{j}P_{ji}/\pi_{i}\times 100\%$. Table I lists these contributions for 2017 and 2019. According to the diagonal numbers, the USA was the most self-sufficient in feeding its competitiveness, with a slight increase from 2017 to 2019. The ninth column also shows that the USA was the most significant contributor to Canada, China, Germany, Japan, and the UK in 2019. For the same year, China was the most significant contributor to Russia’s competitiveness and the second largest to those of Canada, Japan, and the USA. France, Germany, and Italy were highly reliant on each other’s contributions, and their average dependence on Russia was 8.7 times higher than the USA’s dependence on Russia. In value, the USA contributed .14492.025965=.003763.14492.025965.003763.14492*.025965=.003763.14492 ∗ .025965 = .003763 to Canada while Canada contributed .013878.293880=.004078.013878.293880.004078.013878*.293880=.004078.013878 ∗ .293880 = .004078 to the USA in 2019. Comparing the years before and after the beginning of the China-USA trade war, the USA’s contributions to China dropped by $21.7\%$, calculated as 1(1.5849.093905)/(2.1950.086596)11.5849.0939052.1950.0865961-(1.5849*.093905)/(2.1950*.086596)1 - ( 1.5849 ∗ .093905 ) / ( 2.1950 ∗ .086596 ), while China’s contribution to the USA reduced by $15.2\%$, calculated as 1(.8107.29388)/(1.0322.272142)1.8107.293881.0322.2721421-(.8107*.29388)/(1.0322*.272142)1 - ( .8107 ∗ .29388 ) / ( 1.0322 ∗ .272142 ).
@@
Figure VII illustrates the national bargaining power from 2000 to 2019, displaying patterns similar to those in Figure IV, albeit smoother. China’s bargaining power steadily increased due to rapid production growth, even during its $\pi_{i}$’s troughs in 2005 and 2015. Notably, China’s bargaining power consistently surpassed its competitiveness, whereas the USA’s bargaining power was consistently lower than its competitiveness, indicating $\beta$’s mitigation of the Matthew effect in $\pi$. Russia’s bargaining power sharply declined following the Russo-Georgian War in 2008 and the annexation of Crimea in 2014.
Figure VIII illustrates the comparative bargaining power of the USA against other countries. The bargaining power of four Western European countries experienced a slight increase following the establishment of the Eurozone in 1999, but saw a modest decline post-2008, culminating in an average loss of 15.7% of their bargaining power to the USA. Conversely, China and Russia significantly enhanced their relative power multiple times within the first 15 years, maintaining a steady level thereafter. Japan, however, suffered a substantial reduction, losing approximately 50% of its relative bargaining power. Canada’s relative bargaining strength surged during crises but swiftly reverted to its original level, attributable to its strong alignment with the USA.
@@
Table II presents the derivatives of $\log\pi_{i}$ with respect to $\log P_{ji}$ for the years 2000 and 2017, calculated from Eqs. (9) and (12). For instance, a 1% reduction in imports from China by the USA in 2017 would have resulted in an increase of .04889% in the USA’s competitiveness. Although the USA may have had reasons to initiate trade wars against all eight other countries, its primary targets were China in 2017 and Japan in 2000, as indicated by the numerators and denominators in the last row. As evidenced by the last column, China and Japan would have been the two largest beneficiaries had they further collaborated with the USA in 2017 and 2000, respectively. However, a comparison of the numerators and denominators in the column reveals a considerable reduction in benefits from 2000 to 2017. Additionally, initiating trade friction from any of these eight countries against the USA would result in a loss for the initiating country. A trade war against Russia would be nearly fruitless due to the small numerators and denominators in the eighth column. Furthermore, the numbers symmetric with the diagonal exhibit opposite signs, corroborating Corollary 3.
@@
Table III lists the derivatives of $\log\pi$ with respect to $\log P_{ii}$ for 2017 and 2000, calculated from Eqs. (17), (18), and (12). From the diagonal, globalization significantly benefited China, Germany, Italy, Japan, and Russia. It was advantageous for Canada in 2000 but detrimental in 2017. The effects on France were insignificant due to small numerator and denominator. The USA and the UK would be anti-globalization advocates, explaining recent events such as the Trump administration’s withdrawal from international organizations and the UK’s Brexit. Though business interest groups overwhelmingly profited from the gigantic international markets and low-priced raw materials, growing low-income Americans from manufacturing offshoring demanded more inexpensive goods from China, uplifting trade deficits and national debts but benefiting from low inflation and affordable merchandises. Also, China was the largest beneficiary of globalization in 2017 with a 2.324% hike in $\pi_{i}$ for a 1% decrease in $P_{ii}$, so its Belt and Road Initiative (BRI) would have a positive impulse on its $\pi_{i}$. Russia was the largest beneficiary of globalization in 2000.
The side effects on third parties are positioned off the diagonal. For instance, the USA’s further globalization benefited all eight other countries, thereby making its protectionist policies likely to face significant resistance from its partners. China emerged as the largest beneficiary of the USA’s globalization in both years. Conversely, Canada experienced the least impact from the USA’s protectionism in 2017; its further globalization would have marginally benefited the USA, as indicated by the first row. The last column suggests that further globalization from seven other countries would adversely affect the USA; Japan and China topped the list in 2000 and 2017, respectively. Their adverse impact decreased by 70.2% and increased by 380%, respectively, over the 17 years. Additionally, the second row elucidates the USA and its Western allies’ objection to China’s BRI-like projects, which intensified by $4.74$ times from 2000 to 2017, according to the ratios of their numerators and denominators.
@@
We assess the externalities on the third parties due to the USA’s trade war against China in 2018 and 2019. Table IV is derived from Eqs. (9), (11), and (12). When the USA reduced its imports from China by 1% in 2018 and 2019, China’s $\pi_{j}$ suffered a significant loss of .1058% and .0937%, respectively, while the USA’s $\pi_{i}$ earned a non-significant gain of .0540% and .0460%, respectively. Canada was a free rider, and the other six countries were slightly negatively affected in 2018 and 2019. Consequently, these countries would likely refrain from taking formal action to prevent the trade war. Furthermore, Canada or Russia would be the best ally with the USA or China, respectively, in this conflict. The second-best allies would be the UK for the USA and Japan for China, respectively. From 2018 to 2019, the effects were slightly mitigated for all these countries. Likely recognizing the mutually destructive consequences of the trade war, China and the USA agreed on a new trade deal in early 2020 (cf, US Trade Representative, 2020), concluding the first phase of a long-term conflict. However, the cumulative effects would be significant as the conflict is expected to continue for years.
@@
Finally, the cooperative solution $\lambda_{ji}^{*}$ could mitigate the conflict and prevent further escalation. At the inception of this conflict in 2018,
@@
For a positive $\mathrm{d}P_{ji}$, any $\lambda_{ji}$ between $\pi_{j}/\pi_{i}$ and $g_{j}/g_{i}$ would not only make $\frac{\mathrm{d}\log\pi_{i}}{\mathrm{d}\log P_{ji}}<0$ but also create further trade deficits for the USA (see Table V). Thus, an economic conflict with a negative $\mathrm{d}P_{ji}$ was possible and, in this lose-lose conflict, the Nash bargaining solution was:
@@
for 2018. For each dollar’s decrease in exports from China to the USA, $\mathrm{d}P_{ji}=-\$1/g_{j}$. According to the bargaining solution, China would have changed its imports from the USA by $\mathrm{d}P_{ij}=\lambda_{ji}^{*}\mathrm{d}P_{ji}$ of the USA’s production, which is .$66164.49233*(-1/g_{j})*g_{i}=-\$.$66164. Therefore, if this resolution had been applied, the USA would have earned \$.33836=(-.66164)-(-1) surplus for each decreased dollar of imports from China. Additionally, by Theorem  2, the USA’s $\pi_{i}$ would have increased since $\lambda^{*}_{ji}>\pi_{j}/\pi_{i}$ and $\mathrm{d}P_{ji}<0$. To maintain its competitive advantage $\pi_{i}$ and reduce national debt by earning a trade surplus in the bilateral friction, by the impossibility trilemma in Theorem 4 , the USA had to sacrifice its exports and shrink its imports — in either case, $\mathrm{d}P_{ij}<0$.
@@ 7 Discussions
Economic globalization and trade wars have emerged as significant public interests due to global competition among nations, large-scale restructuring of labor markets, and the international distribution of added value. This study posits that import and export data already reflect these issues, and that competitive advantages, rather than comparative advantages, are the primary drivers of trade wars and de-globalization. Consequently, we aim to extract a nation’s competitive strength and bargaining power from the data, identify optimal strategies for addressing trade conflicts and cooperation, and determine the extent to which a nation should pursue globalization.
The methodologies employed, grounded in network and game theories, focus on contemporaneously interactive relationships among economies and are devoid of parameters and pre-set utility functions. These network game approaches can be applied to other bilateral flow or dyadic data (such as traffic, social networks, and global payment systems) and similar conflicts (such as currency wars and technology wars).
Several critical factors warrant further investigation. First, counterbalance disequilibrium may arise when trades are notably imbalanced. This imbalance can manifest in various ways. For instance, when country $i$ uses borrowed money to consume country $j$’s exports, country $i$’s consumption still exerts influence over country $j$’s production. However, this influence originates not only from country $i$, but also from the money lender, who possesses some control over $i$’s consumption. If the lender happens to be country $j$, then $\pi_{i}$ is overestimated and $\pi_{j}$ is underestimated, potentially leading to a debt trap for country $i$.
Secondly, unlike an economic superpower, an emerging economy may prioritize the growth of exports over competitive advantages. Indeed, its economic growth must significantly surpass that of advanced economies to bridge the competitiveness gap, in accordance with $\pi$’s Matthew effect. Thus, maximizing competitive advantages may be a low priority for an emerging economy. For example, during China’s primitive capital accumulation in the 1980s and 1990s, earning additional US dollars was essential to finance its industrial modernization. The time dimension, absent from the current research, is necessary to formulate growth strategies.
Thirdly, the facets of economic globalization encompassing capital, information, and technology have yet to be fully integrated. These elements interact with goods and services, but their data are not as easily accessible. Additionally, several data issues may undermine the empirical study’s results. Notably, there are numerous missing values in the service sectors, and discrepancies often arise between country $i$’s reported imports from country $j$ and country $j$’s reported exports to country $i$.
Furthermore, the integration of regional economies warrants further examination. Theorem 2 could be employed to identify suitable candidates for integration. However, the underlying assumption may need modification. For instance, if country $i$ is a member of multiple RTAs, then $\Delta P_{ii}$ may not be proportionally offset by other countries as used in Theorem 6.
Moreover, additional variables could be incorporated into the demand-sided counterbalance Eq. (3), particularly those from the supply side, such as labor mitigation, which may fine-tune the fractions in $P$. For example, incorporating labor loss into $g_{i}$. Any factors directly influencing country $i$’s production could be included in the $i$th row of $P$. Although consumption may still predominantly drive the fractions, not all production is driven by consumption.
The framework can be extended in several directions. Firstly, amid a bilateral trade war and anti-globalization sentiments, a country may target specific groups of partners without directly impacting others. For instance, as illustrated in Table II, the USA could simultaneously increase tariffs on imports from China, Germany, and Japan.
Secondly, while the dynamical system described by Eqs. (9)–(11) or Eqs. (17)–(18) assumes a fixed $P$, introducing a temporal dimension to $P$ could provide valuable insights. Incorporating a time horizon would facilitate the examination of a trade war’s effects on inflation, $\pi$’s dependence on historical values, and the improvement of trade deficits. Additionally, the temporal dimension could enhance the model’s forecast capabilities. Intertemporal delays often occur when production takes place in one year but consumption in another. If production spans two years, the previous year’s imports of intermediate goods could be included in the current year’s completed production.
Next, we could allow $\lambda_{ji}$ to depend on additional determinants beyond competitiveness and production ratios. An industrial or geopolitical analysis could provide valuable insights into this choice. For instance, as Japan and South Korea competed in automobile and electronic exports in 2000, $\pi_{j}/\pi_{i}$ was appropriate for $\lambda_{ji}$. In the same year, no industry served as a battleground for China and the USA, making $g_{j}/g_{i}$ a natural choice for $\lambda_{ji}$. Moreover, the literature (e.g., Allen, Arkolakis, and Takahashi, 2020; Isard, 1954; Yilmazkuday, 2021) highlights the importance of the distance between trade partners and the type of merchandise as crucial determinants of trade flows. Consequently, one might discount the $i$th row of $P$ by the distances between country $i$ and its trade partners, partially removing location advantage from $\pi$.
Econometric analyses could also provide additional insights. In the linear regression of Eq. (8), for example, if the estimated residual $\hat{\epsilon}_{i}$ is greater or less than 00, we may infer that $\pi_{i}$ is over- or under-valued, respectively. Including additional regressors, such as debt to GDP ratio, in the regression helps to better understand $\pi$ through explanatory data.
Finally, anticipating the side effects of country $i$’s trade war against country $j$, country $j$ might seek cooperation with third parties that would substantially suffer from the conflict. However, adding more countries to both sides could result in a global trade war, akin to the Cold War between 1947 and 1991.
@@
We multiply $\vec{1}_{n}$ from the right on both sides of the equation $\rho P=c\rho$ to get $\rho P\vec{1}_{n}=c\rho\vec{1}_{n}.$ Using $P\vec{1}_{n}=\vec{1}_{n}$ and $\rho\vec{1}_{n}=1$, we obtain $\rho\vec{1}_{n}=c.$ Thus, $c=1$. Also, by $\rho P=c\rho=\rho$ and the uniqueness of $\pi$, we have $\rho=\pi$.
A2. Proof of Corollary 1
Assume that $g_{j}>g_{k}$. For example, country $i$ considers certain imports from either country $j$ or $k$, all else being equal. If $i$ selects $j$, by Eq. (4), then the competitiveness country $i$ derives from $j$ in the imports is approximately
Because of the Matthew effect, ${\pi_{j}}/{g_{j}}>{\pi_{k}}/{g_{k}}$, and so, country $i$ would likely choose $j$, all else being equal.
A3. Proof of Theorem 2
We apply matrix calculus with the restrictions of Eq. (3) and $\pi\vec{1}_{n}=1$ to prove the theorem. When making a small shock or perturbation $\Delta P$ to $P$ in Eq. (3), the new authority distribution $\pi+\Delta\pi$ satisfies the counterbalance equation:
@@
subject to $\Delta P\vec{1}_{n}=\vec{0}_{n}$ and $\Delta\pi\vec{1}_{n}=0$. After subtracting $\pi=\pi P$ from Eq. (A.1), we get $\Delta\pi[I_{n}-P]=\pi\Delta P+\Delta\pi\Delta P$ and its first-order approximation is: